If a person is asked to buy a share of a company where the dividend is expected 7$ next year and he is told that the growth rate is constant at 11% for the foreseeable future what will be the price of the share if the discount rate is 17%?                                                                         


1
Expert's answer
2021-12-14T10:04:14-0500

Given,

Expected dividend, D1 = $7

Growth rate , g = 11%

Discount rate, Ke = 17%

"Price \space of\space share = \frac{D1}{ (Ke - g)}\\

= \frac{\$7 }{ (17\% - 11\%)}\\

= \frac{\$7 }{ 6\%}\\

= \frac{\$7 }{ 0.06}"